OpenStudy (anonymous):

If a triangle has a base of 2 and a half and thhe area is 4 what is the height?

OpenStudy (anonymous):

2.5 * x /2 = 4 2.5x = 8 x = 3.2

OpenStudy (anonymous):

bh/2

OpenStudy (anonymous):

because a triangle is half a rectangle

OpenStudy (anonymous):

b=2.5 a=4 A=bh/2 2A=bh h=2A/b h=2*4/(2.5)

OpenStudy (anonymous):

It's 3.2

OpenStudy (anonymous):

8/2.5

OpenStudy (anonymous):

It works click on the link I submitted